The volcano has continued to be quiet in the last 24 hours, with occasional pyroclastic flows from the south of the dome. Excellent atmospheric conditions have allowed very clear views of the volcano to be obtained.
The volcano has been crystal clear all day. A few small pyroclastic flows were observed from a steep face at the back of the scar above Galway's Wall. As yesterday, the some of the rockfalls and pyroclastic flows began with strong pulses of ash generation, accompanied by the forceful ejection of blocks from the face. There was a lot of gas venting from the northern side of the new extrusion. The wind was generally light and minor amounts of ash from the pyroclastic flows drifted slowly to the west and north.
The level of seismic activity has been low, with only 8 volcano-tectonic earthquakes and 1 long period earthquake. 6 of the volcano-tectonic earthquakes were located at between 1.2 and 2.8 km depth beneath the crater. The rockfall and pyroclastic flow activity has continued at about the same level with 23 signals recorded by the seismic network. Two distant earthquake signals were also recorded by the seismic network.
The long-occupation GPS survey between Harris and the crater rim at Farrells, that was completed yesterday, has been processed today and the results show that the movement of the crater rim away from the dome, that has been seen over the last few measurements, has now stopped.
Total station measurements were made today between Lower and Upper Amersham; the data are currently being processed.
COSPEC measurements were made throughout the day by three methods. In the morning, a normal survey was attempted along the south-west coast road. However, the plume was directed too far to the north to complete a traverse, and so measurements were attempted from the top of Garibaldi Hill. This afternoon, the helicopter was used to attempt traverses under the plume to the north of the volcano. However, atmospheric conditions meant that the plume was spreading in all directions, and so it was difficult to get good quality readings.
Later in the day, temperature measurements were taken at different heights above the volcano from the helicopter. Together with humidity data and the height of the top of the plume, an estimate of the total gas flux from the volcano can be calculated.
The Hot Pond was sampled today by MVO staff and visiting scientists from the Guadeloupe Observatory to analyse the gas and water from the hot springs. This is part of a long term monitoring plan to see how the underground water system may be affected by the volcanic activity.
Observations of the dome made today suggest that major pyroclastic flows to the south could occur at any time and MVO urges everybody to stay alert. Nobody should sleep in zones A to D tonight, which includes Plymouth, St Patrick's, the central corridor, Harris', Long Ground, Spanish Point and Richmond Hill. Zone E, which includes Corkhill and the airport, remains safe at this time.